The measurement of the complement of an angle is 39 degrees more than the angle. Find the measurement of the angle and it complement.
Could you please help me with this problem? thank you so much for your time!

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

The measurement of the complement of an angle is 39 degrees more than the angle. Find the measurement of the angle and it complement.
========================
Call the angle A (in degrees).
The complement is 90 - A.
------
90 - A = A + 39
==========================
My answer is not incorrect.
I don't think tutoring is doing it for you, so I didn't do the arithmetic.
Ikelyn is INCORRECT about that.
She seems to need to feel superior, though she contributes a lot of gibberish.
-------------
All the tutors are here to try to help the students learn. Doing it for them is not the best approach, IMO.
There are a few tutors who seem to like to squabble and complain about the others, for whatever reason.
EG, Ikelyn criticizes the English of the posters, when her English is less than perfect, and never misses an opportunity to say " .... 's answer is INCORRECT."
In CAPS.



Answer by ikleyn(52794)

The measurement of the complement of an angle is 39 degrees more than the angle. Find the measurement of the angle and it complement.
~~~~~~~~~~~~~~~~~~~~~~~


            Since the solution by Alan is INCORRECT,  I came to provide a correct one for you . . .


Let A be the angle you are looking for.


Then the complement is  90 - A degrees.


The condition says that the complement of the angle is 39 degrees more tan the angle. 
It means


    90 - A = A + 39.


    90 - 39 = A + A

    2A = 51  ====>  A = 51/2 = 25.5 degrees = 25 degrees and 30 minutes.

Solved, explained and completed.

-------------------------

To learn the subject, look into the lessons
    - Angles basics
    - Vertical angles
    - HOW TO solve problems on supplementary, complementary or vertical angles - Examples
in this site.